# Physics-Informed Time-Integrated DeepONet: Temporal Tangent Space Operator Learning for High-Accuracy Inference

### v2

@@ 2.3 Reference Methods: Full Rollout and Autoregressive Rollout
To demonstrate the advantages of the proposed approach, we present a comprehensive comparison against state-of-the-art methods, including FR and AR. For FR, the operator is trained on the temporal domain, $t\in[0,0.5]$ for all examples, learning to predict the entire spatiotemporal output field over this interval in a single shot (without respecting causality). In AR, the training process uses three temporal slices per trajectory, specifically at $t\in\{0,0.25,0.5\}$, to learn discrete mappings from each slice to the subsequent step. Within hybrid physics-informed and data-driven formulations, data losses for FR are applied continuously over the temporal domain, while in AR they are localized to the selected time slices with both being supplemented by physics-informed losses. The datasets used for data-driven and physics-informed terms are distinct but consistent across methods, ensuring a fair comparison. This design ensures a consistent and equitable comparison between FR and AR by providing both methods with access to identical training information. Note that the losses here are PDE, IC, BC, and a data loss, where IC refers to the reconstruction loss but is named IC to better reflect its purpose in numerical settings.
@@ 3 Numerical Results
Table 1: Minimum, mean, and maximum relative $\mathcal{L}_{2}$ error along with mean and standard deviation for 5 timed runs each over all test examples for three test cases in comparison of FR, AR, and time integration using explicit Euler, fourth-order Runge-Kutta (RK4), and second-order Adam-Bashforth-Moultons (ABM2) as inference scheme with $\Delta t=0.01$. Irrespective of the inference scheme employed, PITI‑DeepONet decreases the mean relative $\mathcal{L}_{2}$ errors by 84% (vs. FR) and 79% (vs. AR) for the one‑dimensional heat equation; by 87% (vs. FR) and 98% (vs. AR) for the one‑dimensional Burgers’ equation; and by 42% (vs. FR) and 89% (vs. AR) for the two‑dimensional Allen–Cahn equation. The timing includes residual calculation for PITI cases.
Heat 1D @ $t=5.0$ Burgers 1D @ $t=1.0$ Allen-Cahn 2D @ $t=1.0$
Min / Mean / Max Inference in $s$ Min / Mean / Max Inference in $s$ Min / Mean / Max Inference in $s$
FR 0.82 / 1.5 / 24 $1.4\mathrm{e}\text{-}3\!\pm\!1.7\mathrm{e}\text{-}5$ 0.048 / 0.13 / 0.56 $5.8\mathrm{e}\text{-}4\!\pm\!3.1\mathrm{e}\text{-}5$ 0.13 / 0.20 / 0.28 $9.5\mathrm{e}\text{-}4\!\pm\!1.7\mathrm{e}\text{-}5$
AR 0.11 / 1.2 / 23 $6.8\mathrm{e}\text{-}2\!\pm\!3.9\mathrm{e}\text{-}5$ 0.15 / 0.94 / 11 $1.6\mathrm{e}\text{-}3\!\pm\!7.6\mathrm{e}\text{-}5$ 0.49 / 1.1 / 1.7 $9.2\mathrm{e}\text{-}3\!\pm\!2.2\mathrm{e}\text{-}4$
PITI (Euler) 0.029 / 0.24 / 3.4 $7.5\mathrm{e}\text{-}2\!\pm\!1.7\mathrm{e}\text{-}5$ 0.0084 / 0.018 / 0.26 $1.4\mathrm{e}\text{-}2\!\pm\!9.0\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $8.9\mathrm{e}\text{-}3\!\pm\!3.1\mathrm{e}\text{-}5$
PITI (RK4) 0.029 / 0.24 / 3.3 $2.0\mathrm{e}\text{-}1\!\pm\!2.6\mathrm{e}\text{-}4$ 0.0038 / 0.017 / 0.26 $3.8\mathrm{e}\text{-}2\!\pm\!8.5\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $2.6\mathrm{e}\text{-}2\!\pm\!3.7\mathrm{e}\text{-}5$
PITI (ABM2) 0.029 / 0.24 / 3.3 $1.7\mathrm{e}\text{-}2\!\pm\!2.2\mathrm{e}\text{-}4$ 0.0039 / 0.017 / 0.26 $3.2\mathrm{e}\text{-}2\!\pm\!3.6\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $2.4\mathrm{e}\text{-}3\!\pm\!6.1\mathrm{e}\text{-}5$
@@
Figure 2: Relative $\mathcal{L}_{2}$ error over time for the three PDEs in comparison of FR, AR, and time integration using explicit Euler with $\Delta t=0.01$. Depicted are the mean as well as minimum and maximum value for all test examples. Only inference with explicit Euler is depicted, but all methods show similar values in inference also as shown in Tab. 1.
@@
The subsequent sections detail the setup and results for the three examples of Heat and Burgers’ equations in one dimension and Allen-Cahn equation in two dimensions. Fig. 2 and Tab. 1 present a short overview of the performance of our proposed method compared to FR and AR. The relative $\mathcal{L}_{2}$ error = $\|\hat{\mathbf{u}}^{n}-\mathbf{u}^{n}\|_{2}/\|\mathbf{u}^{n}\|_{2}$ was employed on the predicted field as standard metric for comparison with the field, depicting relative deviation from the ground truth. Furthermore, all hyperparameters were optimized using optuna [akiba2019optuna]. Metric for AR and FR models was to reduce the relative $\mathcal{L}_{2}$ error on the complete validation data for the field output, while PITI models were optimized on the relative $\mathcal{L}_{2}$ error of the time derivative within the complete validation data. Optimization was performed on a discrete set of pre-chosen options to ease interpretation. At least 20 runs with a minimum of 6h run time were performed per setup. This included depth and width of branch and trunk nets, hidden dimension, training epochs, learning rate settings, loss weights, as well as information on how to obtained multiple outputs by splitting either or both trunk and branch outputs where necessary. For networks with CNNs in the branch, the number of layers, kernels per layer, depth and size of the dense layers, if pooling was applied and if so whether average or max pooling, as well as dense and CNN activation. Stride for convolutions and pooling as well as kernel size for convolutions were fixed to 2 in each dimension. Interestingly, all CNNs converged to the same structure including activation functions except for width of the dense layer. The results are given in Tab. 2 for everything except loss weights which are presented in Tab. 3. Random seed choice in physics-informed scientific machine learning can influence solution diversity and optimization outcomes, especially in problems with multiple valid solutions, since random initialization may lead to discovering different solutions [zou2025learningdiscoveringmultiplesolutions]. However, as the obtained results showed good agreement during training with the used input, a single random seed per numerical examples was utilized. This is a standard, resource-efficient practice within the field, but we acknowledge that additional seeds might yield alternative results in non-convex or multi-solution settings. Training and inference were performed on an NVIDIA A40 GPU with 48 GB of memory, utilizing CUDA version 12.4 and NVIDIA driver version 550.54.15, with the GPU operating in a Linux environment running the 64-bit Debian-based GNU/Linux operating system under kernel version 6.1.0-34-cloud-amd64. The framework was built in Python 3.11.2 based on JAX [jax2018github] and Flax [flax2020github] with integration of tools from the DeepMind Ecosystem [deepmind2020jax]. The full code included utilized package versions will be available at https://github.com/lmandl/PITI-DeepONet upon publication of this work. The dataset for the underlying cases was generated from scratch for this paper. While some datasets exists and either the data or the code to generate the dataset are publicly available (cf. [Wang2021] for Burger’s equation), so far no dataset includes high-accuracy time derivatives. Hence, the need for this data to track accuracy of our method or additional input for hybrid data-driven and physics-informed approaches, required generation of new datasets that fit the purpose of solving PDEs.
@@
Table 2: Hyperparameter values after optuna optimization for all setups. Branch inputs are 128 sensors for the heat equation, 101 sensors for Burgers’ equation, and $(16,16)$ for the Allen-Cahn equation. Trunk input dimensions are 2 for the heat and Burgers’ equation $(t,x)$ and 3 for the Allen-Cahn equation $(t,x,y)$. For models with the special form for time-evolution PDEs, the trunk input is reduced by the time dimension. p denotes latent dimension for the branch and trunk nets. All models are unstacked DeepONet that are optimized with Adam optimizer and exponential learning rate decay. Hy. indicates a hybrid data-driven and physics-informed model, while spec. indicates the special form. The architecture of the convolutional neural network (CNN) used for all three Allen-Cahn (AC) cases consists of an initial 2D convolution (Conv) with a stride of 2 and 2 filters, followed by average pooling (AvgPool) with a stride of 2, then another 2D convolution with a stride of 2 and 4 filters, and another average pooling with a stride of 2. The output is then flattened and passed through a dense layer, with the size specified in the column for the branch network. Activation function for all models including the CNNs is hyperbolic tangent. Timing is averaged over the whole training including 100 validation evaluations equally distributed throughout training.
@@
Table 3: Loss weights after optuna optimization for all setups. Factors for reconstruction loss in time integration and initial condition loss in rollout settings are combined into one column as they essentially have the same significance.
@@ 3.1 One-dimensional Heat Equation
As an introductory example, we present the 1D heat equation with the form:
@@
The dataset is generated by solving Eq. 9 with Dirichlet BCs on on the mentioned spatiotemporal domain. A uniform grid with $N_{x}=128$ points is used for space, and the solution is computed over $N_{t,\mathrm{fine}}=10,000$ fine time steps with a step size $\Delta t_{\mathrm{fine}}$, which inherently satisfies the stability requirements of the numerical scheme. The IC $u_{0}(x)$ is sampled from a periodic Gaussian random field. The numerical solution is obtained using an explicit finite-difference scheme, with second-order accurate spatial derivatives in the interior and fourth-order one-sided differences at the boundaries. The time derivative $\partial_{t}u$ is computed at each step, and BCs are enforced by setting $u(t,0)=u(t,1)=0$. The solution is then downsampled to a coarse temporal grid with $N_{t}=100$ intervals using linear interpolation, yielding the coarse fields $u(t,x)$ and $\partial_{t}u(t,x)$. The dataset includes $N_{\mathrm{train}}=2,000$ samples for training and $N_{\mathrm{test}}=500$ for testing, with solutions recorded for $T=1$ and $T_{\mathrm{test}}=5$, respectively.
@@
Final mean relative $\mathcal{L}_{2}$ errors observed were $6.81\mathrm{e}\text{-}2$ on the training set and $8.20\mathrm{e}\text{-}2$ on the validation set for the time derivative using TI, whereas for the field values, AR achieved $4.09\mathrm{e}\text{-}3$ and $2.04\mathrm{e}\text{-}3$, and FR achieved $6.63\mathrm{e}\text{-}3$ and $7.20\mathrm{e}\text{-}3$ on the training and test sets, respectively (see Fig. 4 for full trajectories). Although FR and AR achieve an order of magnitude better results on the training domain $t\in[0,1]$, if used in inference, PITI beats AR on $t\in[0,1]$ and both AR and FR on $t\in[0,5]$ as shown in Fig. 2(a) and Tab. 1 with full field plots and differences for a single test example for PITI using explicit Euler, AR, and FR in Fig. 7. Furthermore, all employed time integration methods show near identical results as per Tab. 1 as well as Fig. 6 which depicts behavior over time including residuals. Fig. 3 illustrates inference using explicit Euler, alongside the squared difference between the reference and prediction, as well as the calculated residual on the network’s output. As hypothesized, the squared difference and the residual exhibit similar patterns, highlighting the potential of using the residual as an indicator of prediction quality. This relationship becomes even more pronounced when computing the mean absolute error (MAE) and the Pearson correlation coefficient $\rho$ over all test examples, yielding a MAE of $2.04\mathrm{e}\text{-}3$ and $\rho=0.997$, thereby indicating an almost perfect correlation across the dataset.
@@
As mentioned in Section 2.2.1, a special formulation can be employed for time-evolution PDEs, such as the heat equation, where the time derivative appears in isolation and all other terms can be moved to the right-hand side. This alternative setup demonstrates comparable performance to the standard (vanilla) formulation, with mean relative $\mathcal{L}_{2}$ errors of $2.39\mathrm{e}\text{-}1$ for the standard and $2.48\mathrm{e}\text{-}1$ for the special formulation (Fig. 5).
@@
Figure 5: 1D Heat equation: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit Euler scheme while depicting minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
Figure 6: 1D Heat equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
Figure 7: 1D Heat equation: Comparison of predictions made using time integration (PITI) with explicit Euler, autoregressive and full rollout alongside the differences for each method.

### v3

@@ 2.3 Reference Methods: Full Rollout and Autoregressive Rollout
To demonstrate the advantages of the proposed approach, we present a comprehensive comparison against state-of-the-art methods, including FR and AR. For FR, the operator is trained on the temporal domain, $t\in[0,0.5]$ for all examples, learning to predict the entire spatiotemporal output field over this interval in a single shot (without respecting causality). In AR, the training process uses three temporal slices per trajectory, specifically at $t\in\{0,0.25,0.5\}$, to learn discrete mappings from each slice to the subsequent step. Within hybrid physics-informed and data-driven formulations, data losses for FR are applied continuously over the temporal domain, while in AR, they are localized to the selected time slices, with both being supplemented by physics-informed losses. The physics-informed collocation is paradigm-dependent: FR enforces the residual over the full training horizon $t\in[0,T]$, whereas AR enforces it only over a single-step window starting from the current time $t^{n}$ (equivalently acting on a local interval of length $\Delta t$, i.e., $0,\Delta t]$, per transition), and PITI enforces it at the current-time mapping only (equivalently at a single time point, i.e., $[0]$). In hybrid settings, the supervised targets are likewise method-specific: FR uses trajectory snapshots, AR uses state pairs $(\mathbf{u}^{n},\mathbf{u}^{n+1})$, and PITI uses time-derivative targets $\mathbf{u}_{t}^{n}$, but in all cases they are derived from the same reference trajectories for identical initial conditions, ensuring a fair comparison. This design ensures a consistent and equitable comparison between FR and AR by providing both methods with access to identical training information. Note that the losses here are PDE, IC, BC, and a data loss, where IC refers to the reconstruction loss but is named IC to better reflect its purpose in numerical settings. Beyond FR and AR, we incorporate two additional baselines from the literature for 1D Burgers’ equation. First, we consider physics-constrained deep autoregressive networks (PCAR; Geneva and Zabaras [2020]), which advance the solution by repeatedly applying a state-to-state convolutional predictor that maps the current (and optionally several past) spatial fields to the next field with a fixed time step. Architecturally, this predictor is a dense encoder–decoder convolutional neural network consisting of encoding convolutions, a dense block, and a decoding block that upsamples to the original grid. Training is purely physics-constrained by minimizing a discretized time-integrator residual (e.g., Crank–Nicolson / Euler variants), with spatial derivatives approximated via finite-difference-like stencil convolutions, thereby eliminating the need for labeled trajectory pairs or automatic differentiation. To obtain a purely physics-informed PCAR baseline under the same training data and hyperparameter budget as our AR and PITI variants, we retrained PCAR on our dataset and adapted it to our setup by using a single input slice instead of multiple previous slices. Second, we include the physics-informed neural ODE (PINODE; Sholokhov et al. [2023]), a reduced-order modeling framework that jointly learns an autoencoder (encoder–decoder pair) and a neural model of time evolution in a low-dimensional latent space. Specifically, the autoencoder compresses the physical state into latent variables, a neural network governs their temporal evolution, and an ODE solver integrates this evolution forward in time before decoding back to the original state. Its physics-informed training enforces consistency with the governing dynamics via collocation: the learned latent evolution is regularized so that, when passed through the encoder/decoder, it matches the known physical evolution while maintaining accurate reconstruction, which necessitates a differentiable encoder and numerically stable automatic differentiation. In contrast to discrete autoregressive rollouts, PINODE operates intrinsically in continuous time, so its long-horizon behavior is primarily determined by the latent dimensionality, the expressiveness of the latent dynamics network, and the numerical stability of the ODE integration, rather than by the number of rollout steps used in training. For PINODE, we report only the published results for the 1D Burgers’ equation (rather than re-implementing the method, as no code was provided alongside the work) and use these as an external reference for long-horizon rollout accuracy in the literature.
@@ 3 Numerical Results
Table 1: Minimum, mean, and maximum relative $\mathcal{L}_{2}$ error along with mean and standard deviation for 5 timed runs each over all test examples for Heat, Burgers’, Allen-Cahn, and Kuramoto-Sivashinsky (KS) equation in comparison of FR, AR, and time integration using explicit Euler, fourth-order Runge-Kutta (RK4), second-order Adam-Bashforth-Moultons (ABM2) and implicit Euler (IE) as inference scheme with $\Delta t=0.01$. Irrespective of the inference scheme employed, PITI‑DeepONet decreases the mean relative $\mathcal{L}_{2}$ errors by 84% (vs. FR) and 79% (vs. AR) for the one‑dimensional heat equation; by 87% (vs. FR) and 98% (vs. AR) for the one‑dimensional Burgers’ equation; by 42% (vs. FR) and 89% (vs. AR) for the two‑dimensional Allen–Cahn equation; and by 58% (vs. FR) and 61% (vs. AR) for the one-dimensional Kuramoto-Sivashinsky equation. The timing includes residual calculation for PITI cases.
Heat 1D @ $t=5.0$ Burgers 1D @ $t=1.0$ Allen-Cahn 2D @ $t=1.0$ KS 1D @ $t=10.0$
Min / Mean / Max Inference in $s$ Min / Mean / Max Inference in $s$ Min / Mean / Max Inference in $s$ Min / Mean / Max Inference in $s$
FR 0.82 / 1.5 / 24 $1.4\mathrm{e}\text{-}3\!\pm\!1.7\mathrm{e}\text{-}5$ 0.048 / 0.13 / 0.56 $5.8\mathrm{e}\text{-}4\!\pm\!3.1\mathrm{e}\text{-}5$ 0.13 / 0.20 / 0.28 $9.5\mathrm{e}\text{-}4\!\pm\!1.7\mathrm{e}\text{-}5$ 0.71 / 1.11 / 2.55 $8.6\mathrm{e}\text{-}4\!\pm\!2.6\mathrm{e}\text{-}5$
AR 0.11 / 1.2 / 23 $6.8\mathrm{e}\text{-}2\!\pm\!3.9\mathrm{e}\text{-}5$ 0.15 / 0.94 / 11 $1.6\mathrm{e}\text{-}3\!\pm\!7.6\mathrm{e}\text{-}5$ 0.49 / 1.1 / 1.7 $9.2\mathrm{e}\text{-}3\!\pm\!2.2\mathrm{e}\text{-}4$ 0.27 / 1.20 / 2.17 $7.0\mathrm{e}\text{-}3\pm 3.2\mathrm{e}\text{-}5$
PITI Euler 0.029 / 0.24 / 3.4 $7.5\mathrm{e}\text{-}2\!\pm\!1.7\mathrm{e}\text{-}5$ 0.0084 / 0.018 / 0.26 $1.4\mathrm{e}\text{-}2\!\pm\!9.0\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $8.9\mathrm{e}\text{-}3\!\pm\!3.1\mathrm{e}\text{-}5$ 0.074 / 0.46 / 1.36 $1.4\mathrm{e}\text{-}2\pm 2.6\mathrm{e}\text{-}5$
RK4 0.029 / 0.24 / 3.3 $2.0\mathrm{e}\text{-}1\!\pm\!2.6\mathrm{e}\text{-}4$ 0.0038 / 0.017 / 0.26 $3.8\mathrm{e}\text{-}2\!\pm\!8.5\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $2.6\mathrm{e}\text{-}2\!\pm\!3.7\mathrm{e}\text{-}5$ - -
ABM2 0.029 / 0.24 / 3.3 $1.7\mathrm{e}\text{-}2\!\pm\!2.2\mathrm{e}\text{-}4$ 0.0039 / 0.017 / 0.26 $3.2\mathrm{e}\text{-}2\!\pm\!3.6\mathrm{e}\text{-}5$ 0.079 / 0.12 / 0.16 $2.4\mathrm{e}\text{-}3\!\pm\!6.1\mathrm{e}\text{-}5$ - -
IE - - - - - - 0.083 / 0.47 / 1.42 $1.4\pm 7.7\mathrm{e}\text{-}2$
@@
Figure 2: Relative $\mathcal{L}_{2}$ error over time for the four PDEs in comparison of FR, AR, and time integration using explicit Euler with $\Delta t=0.01$. Depicted are the mean, as well as the minimum and the maximum values, for all test examples. Only inference with explicit Euler is depicted, but all methods show similar values in inference, also as shown in Tab. 1.
@@
The subsequent sections detail the setup and results for the examples of Heat and Burgers’ equations in one dimension and Allen-Cahn equation in two dimensions and close with Kuramoto–Sivashinsky (KS) equation in 1D as an example for chaotic behaviour. Fig. 2 and Tab. 1 present a short overview of the performance of our proposed method compared to FR and AR. The relative $\mathcal{L}_{2}$ error = $\|\hat{\mathbf{u}}^{n}-\mathbf{u}^{n}\|_{2}/\|\mathbf{u}^{n}\|_{2}$ was employed on the predicted field as standard metric for comparison with the field, depicting relative deviation from the ground truth. Furthermore, all hyperparameters were optimized using optuna Akiba et al. [2019]. The metric for AR and FR models was to reduce the relative $\mathcal{L}_{2}$ error on the complete validation data for the field output, while PITI models were optimized on the relative $\mathcal{L}_{2}$ error of the time derivative within the complete validation data. Optimization was performed on a discrete set of pre-chosen options to ease interpretation. At least 20 runs with a minimum of 6h run time were performed per setup. This included depth and width of branch and trunk nets, hidden dimension, training epochs, learning rate settings, loss weights, as well as information on how to obtain multiple outputs by splitting either or both trunk and branch outputs where necessary. For networks with CNNs in the branch, the number of layers, kernels per layer, depth, and size of the dense layers, if pooling was applied, and if so, whether average or max pooling, as well as dense and CNN activation. Stride for convolutions and pooling, as well as kernel size for convolutions, were fixed to 2 in each dimension. Interestingly, all CNNs converged to the same structure, including activation functions, except for the width of the dense layer. The results are given in Tab. 2 for everything except loss weights, which are presented in Tab. 3. Random seed choice in physics-informed scientific machine learning can influence solution diversity and optimization outcomes, especially in problems with multiple valid solutions, since random initialization may lead to discovering different solutions Zou et al. [2025]. However, as the obtained results showed good agreement during training with the used input, a single random seed per numerical example was utilized. This is a standard, resource-efficient practice within the field, but we acknowledge that additional seeds might yield alternative results in non-convex or multi-solution settings. Training and inference were performed on an NVIDIA A40 GPU with 48 GB of memory, utilizing CUDA version 12.4 and NVIDIA driver version 550.54.15, with the GPU operating in a Linux environment running the 64-bit Debian-based GNU/Linux operating system under kernel version 6.1.0-34-cloud-amd64. The framework was built in Python 3.11.2 based on JAX Bradbury et al. [2018] and Flax Heek et al. [2023] with integration of tools from the DeepMind Ecosystem DeepMind et al. [2020]. The full code included utilized package versions will be available at https://github.com/lmandl/PITI-DeepONet upon publication of this work. The dataset for the underlying cases was generated from scratch for this paper. While some datasets exist and either the data or the code to generate the dataset are publicly available (cf. Wang et al. [2021] for Burger’s equation), so far no dataset includes high-accuracy time derivatives. Hence, the need for this data to track the accuracy of our method or additional input for hybrid data-driven and physics-informed approaches required the generation of new datasets that fit the purpose of solving PDEs.
@@
Table 2: Hyperparameter values after optuna optimization for all setups. Branch inputs are 128 sensors for the heat equation and Kuramoto-Sivashinsky (KS) equation, 101 sensors for Burgers’ equation, and $(16,16)$ for the Allen-Cahn equation. Trunk input dimensions are 2 for the heat, Burgers’ and KS equation $(t,x)$ and 3 for the Allen-Cahn equation $(t,x,y)$. For models with the special form for time-evolution PDEs, the trunk input is reduced by the time dimension. p denotes the latent dimension for the branch and trunk nets. All models are unstacked DeepONet models that are optimized with an Adam optimizer and exponential learning rate decay. Hy. indicates a hybrid data-driven and physics-informed model, while spec. indicates the special form. The architecture of the convolutional neural network (CNN) used for all three Allen-Cahn (AC) cases consists of an initial 2D convolution (Conv) with a stride of 2 and 2 filters, followed by average pooling (AvgPool) with a stride of 2, then another 2D convolution with a stride of 2 and 4 filters, and another average pooling with a stride of 2. The output is then flattened and passed through a dense layer, with the size specified in the column for the branch network. Activation function for all models, including the CNNs, is the hyperbolic tangent, unless stated otherwise. Timing is averaged over the whole training, including 100 validation evaluations equally distributed throughout training.
@@
KS PITI Hy. 10x[64] 6x[256] 256 200,000 $1\mathrm{e}\text{-}4$ 0.85 30,000 False True 259.77
AR Hy. 4x[64] 4x[64] 128 200,000 $5\mathrm{e}\text{-}4$ 0.8 20,000 False False 564.62
FR Hy. 10x[64] 4x[256] 256 50,000 $5\mathrm{e}\text{-}4$ 0.9 50,000 False False 23.83
@@
Sine as trunk activation function
Table 3: Loss weights after optuna optimization for all setups. Factors for reconstruction loss in time integration and initial condition loss in rollout settings are combined into one column, as they essentially have the same significance.
@@ 3.1 One-dimensional Heat Equation
As an introductory example, we present the 1D heat equation in the form:
@@
The dataset is generated by solving Eq. 12 with Dirichlet BCs on the mentioned spatiotemporal domain. A uniform grid with $N_{x}=128$ points is used for space, and the solution is computed over $N_{t,\mathrm{fine}}=10,000$ fine time steps with a step size $\Delta t_{\mathrm{fine}}$, which inherently satisfies the stability requirements of the numerical scheme. The IC $u_{0}(x)$ is sampled from a periodic Gaussian random field. The numerical solution is obtained using an explicit finite-difference scheme, with second-order accurate spatial derivatives in the interior and fourth-order one-sided differences at the boundaries. The time derivative $\partial_{t}u$ is computed at each step, and BCs are enforced by setting $u(t,0)=u(t,1)=0$. The solution is then downsampled to a coarse temporal grid with $N_{t}=100$ intervals using linear interpolation, yielding the coarse fields $u(t,x)$ and $\partial_{t}u(t,x)$. The dataset includes $N_{\mathrm{train}}=2,000$ samples for training and $N_{\mathrm{test}}=500$ for testing, with solutions recorded for $T=1$ and $T_{\mathrm{test}}=5$, respectively.
@@
Final mean relative $\mathcal{L}_{2}$ errors observed were $6.81\mathrm{e}\text{-}2$ on the training set and $8.20\mathrm{e}\text{-}2$ on the validation set for the time derivative using TI, whereas for the field values, AR achieved $4.09\mathrm{e}\text{-}3$ and $2.04\mathrm{e}\text{-}3$, and FR achieved $6.63\mathrm{e}\text{-}3$ and $7.20\mathrm{e}\text{-}3$ on the training and test sets, respectively (see Fig. 4 for full trajectories). Although FR and AR achieve an order of magnitude better results on the training domain $t\in[0,1]$, if used in inference, PITI beats AR on $t\in[0,1]$ and both AR and FR on $t\in[0,5]$ as shown in Fig. 2(a) and Tab. 1 with full field plots and differences for a single test example for PITI using explicit Euler, AR, and FR in Fig. 7. Furthermore, all employed time integration methods show near identical results as per Tab. 1 as well as Fig. 6, which depicts behavior over time, including residuals. Fig. 3 illustrates inference using explicit Euler, alongside the squared difference between the reference and prediction, as well as the calculated residual on the network’s output. As hypothesized, the squared difference and the residual exhibit similar patterns, highlighting the potential of using the residual as an indicator of prediction quality. This relationship becomes even more pronounced when computing the mean absolute error (MAE) and the Pearson correlation coefficient $\rho$ over all test examples, yielding a MAE of $2.04\mathrm{e}\text{-}3$ and $\rho=0.997$, thereby indicating an almost perfect correlation across the dataset.
@@
As mentioned in Section 2.2.1, a special formulation can be employed for time-evolution PDEs, such as the heat equation, where the time derivative appears in isolation, and all other terms can be moved to the right-hand side. In our experiments, this efficiency-oriented variant yields comparable accuracy to the standard (vanilla) formulation, with mean relative $\mathcal{L}_{2}$ errors of $2.39\mathrm{e}\text{-}1$ for the standard and $2.48\mathrm{e}\text{-}1$ for the special formulation (Fig. 5).
@@
Figure 5: 1D Heat equation: Comparison of time integration models in the vanilla form and special form for time-evolution partial differential equations, including total loss and mean relative $\mathcal{L}_{2}$ error on train and validation set during training. Only the relative $\mathcal{L}_{2}$ error on the time derivative is provided here. The right plot shows the performance during inference with the fully trained model using an explicit Euler scheme while depicting the minimum, mean, and maximum value of the relative $\mathcal{L}_{2}$ error per time step.
Figure 6: 1D Heat equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples. Note that in both subplots several curves overlap almost exactly and therefore appear as a single trace.
Figure 7: 1D Heat equation: Comparison of predictions made using time integration (PITI) with explicit Euler, autoregressive, and full rollout alongside the differences for each method.